# The Merging Galaxy Cluster Environment Affects the Morphology of Radio-AGN

### v1

@@ Abstract
It has previously been found that the galaxy cluster environment can affect the fueling and evolution of Active Galactic Nuclei (AGN). This work examines the effect of the merging cluster environment on the properties of radio-AGN by comparing the radio morphology of cluster members in a sample of four merging and eight relaxed galaxy clusters at low redshift (z0.2). Using 144-MHz data from the LOFAR Two-metre Sky Survey (LoTSS) and Zooniverse, we classify the radio morphology of the radio-detected cluster members using the following morphology classes: compact, compact extended, extended, jetted, and disturbed. We find that the merging cluster environment has a statistically significant, higher population proportion of disturbed (bent and head tail) sources, indicating that the merging environment can affect the morphology of cluster radio-AGN. We also investigate the number of AGN that are detected in the radio data only, and the number that are detected in both the radio and optical data in mergers and non-mergers. We find that the merging cluster environment has a higher population proportion of AGN that are identified only as radio-AGN compared to AGN that are identified as both radio and optical AGN. Overall, we find that the merging environment affects certain radio-AGN (disturbed and only radio identified AGN), but not all.
@@ 1 Introduction
It is now well established that dense environments such as galaxy clusters affect the properties and evolution of the galaxies within them (see references detailed below). There are a variety of interactions between the galaxies, the intra-cluster medium (ICM; a diffuse, hot, ionized plasma that is gravitationally bound to the cluster), and the cluster potential that are possible and prevalent in clusters that affect properties and evolution of galaxies. Some examples of the possible interactions are galaxy-galaxy harassment (repeated close encounters with neighbors; Moore et al. 1996, 1999), galaxy-cluster potential tidal effects (Fujita, 1998; Natarajan et al., 1998), and galaxy interactions with the ICM, such as ram pressure stripping (Gunn & Gott, 1972) and strangulation (Larson et al., 1980). These interactions can affect the star formation rates (Kauffmann et al., 2004; Fassbender et al., 2014; Barsanti et al., 2018), active galactic nuclei (AGN) activity (Galametz et al., 2009; Bufanda et al., 2017; Poggianti et al., 2017; Marshall et al., 2018; Mo et al., 2018), morphology (Dressler, 1980; Kauffmann et al., 2004; Sazonova et al., 2020), and gas supply of the cluster galaxies (Stroe et al., 2015; Poggianti et al., 2017; Jaffé et al., 2018; Cairns et al., 2019).
@@
One classic example of the effect of the cluster environment on its galaxies is the morphology–density relation, in which, with increasing density, the fraction of elliptical galaxies increases and the fraction of spiral galaxies, decreases such that clusters typically have an overabundance of elliptical galaxies compared to the field (Dressler, 1980; Kodama et al., 2001; Kauffmann et al., 2004; Sazonova et al., 2020). Coupled to the morphology-density relation is the star formation rate, where more star-formation is seen in galaxies at low densities and less at high densities (Kauffmann et al., 2004; Fassbender et al., 2014). There is an evolution of these properties as a function of redshift, in that, with decreasing redshift, the fraction of “red and dead” galaxies increases (De Lucia et al., 2004; Rudnick et al., 2009) and the fraction of star forming galaxies decreases (Margoniner et al., 2001; Poggianti et al., 2006; Saintonge et al., 2008). Another example of the environment affecting cluster galaxies is the case of “jellyfish” galaxies, which are galaxies typically located near the cores of the clusters that are experiencing ram pressure stripping, leaving a tail of gas behind them opposite to the direction of motion (Fumagalli et al., 2014; Ebeling et al., 2014; Jaffé et al., 2018).
@@
The cluster environment is also known to affect the properties and evolution of cluster galaxies that host AGN. AGN are powerful astrophysical sources that are fueled by accretion of matter onto the supermassive black hole at the center of the galaxy and emit characteristic radiation signatures across the full electromagnetic spectrum (Padovani et al., 2017). AGN can be identified in one or multiple wavelengths such as X-ray, optical, infrared, or radio. At all redshifts Galametz et al. (2009); Mo et al. (2018) find that there is an excess of radio identified AGN in clusters, and specifically at the center of clusters compared to the field, which is more pronounced at high redshift (Galametz et al., 2009; Mo et al., 2020). Similarly, pertaining to X-ray, optical, and infrared identified AGN, the number density of AGN continually increases as a function of redshift in that there is a deficit of these types of AGN near the cluster center at (Pimbblet et al., 2013; Ehlert et al., 2014; Gordon et al., 2018), but at there is an excess of these types of AGN in clusters (Ruderman & Ebeling, 2005; Galametz et al., 2009; Fassbender et al., 2012; Martini et al., 2013; Alberts et al., 2016; Bufanda et al., 2017) which is most pronounced near the cluster center (Ruderman & Ebeling, 2005; Galametz et al., 2009; Fassbender et al., 2012; Alberts et al., 2016). The cluster environment seems to affect many different properties of radio-AGN in particular, such as the size, luminosity, and morphology of the radio-AGN (Prestage & Peacock, 1988; Gendre et al., 2013; Ineson et al., 2015; Croston et al., 2019; Garon et al., 2019; Moravec et al., 2020; Macconi et al., 2020; Shen et al., 2020; Morris et al., 2022). Further, bent-tail radio galaxies are a very well-known type of radio-AGN that showcase the effect of the environment on radio-AGN produced by the ram pressure of the dense environment (Miley et al., 1972; Begelman et al., 1979; O’Donoghue et al., 1993; Hardcastle et al., 2005; Morsony et al., 2013).
@@
Galaxy clusters undergo cluster-cluster mergers (henceforth mergers111In this work we refer to cluster-cluster mergers as “mergers” which we differentiate from the typical galaxy-galaxy mergers that many in the literature refer to as “mergers.”) as a result of hierarchical structure formation in the Universe (Voit, 2005; Springel et al., 2005; Boylan-Kolchin et al., 2009; Klypin et al., 2011; Pillepich et al., 2018; Planck Collaboration et al., 2020). Mergers have an immense and obvious impact on the properties of the ICM. For example mergers cause the ICM distribution to be asymmetric seen by X-ray (Buote, 2002) and Sunyaev–Zeldovich observations (Mroczkowski et al., 2019). Further, mergers drive intergalactic shocks and induce turbulence in the ICM (Sarazin, 2002; Markevitch, 2006; Basu et al., 2016). And lastly mergers can create cold fronts in the ICM (Sarazin, 2002; Owers et al., 2009; Roediger et al., 2013) and disrupt the cool cores of clusters (Sarazin, 2002).
@@
Although the effect of mergers on the ICM is clear, the impact on the cluster galaxies is less so. There is contradictory evidence that mergers can either increase (Owen et al., 1999; Miller & Owen, 2003; Hwang & Lee, 2009) or decrease (Hwang & Lee, 2009; Mansheim et al., 2017) star formation in cluster galaxies. One study found that a dynamically younger cluster had a higher fraction of galaxies with a high star formation rate (0.17 M⊙ yr-1, Stroe et al. 2015). There is also evidence that a brief period of enhanced, triggered star formation may be followed by subsequent quenching (Stroe et al., 2015; Cairns et al., 2019). These results are attributed to a disturbed ICM affecting the molecular gas reservoir of the cluster galaxies, with the disturbance exciting star formation and leading to a rapid consumption of the molecular gas (Stroe et al., 2015; Cairns et al., 2019).
@@
Recent work has found that the AGN population is impacted by cluster-cluster mergers as well. The radio-AGN population in particular has been studied in depth in individual merging systems and found to have a higher frequency compared to The radio-AGN frequency in more relaxed systems (e.g., Miller & Owen, 2003; Moravec et al., 2020). Similarly, Noordeh et al. (2020) found enhanced X-ray AGN activity in the most dynamically disturbed of the seven clusters they studied. Bilton et al. (2020) find that mergers hold relatively kinematically younger AGN sub-populations that have recently coalesced into a common potential (similar to a first infall population of galaxies) compared to relaxed clusters. In general, Miller & Owen (2003); Hwang & Lee (2009); Sobral et al. (2015) find that mergers can provoke AGN activity. However, the physical mechanism by which mergers instigate AGN activity and its effects on radio-AGN properties is unknown.
@@
In this paper, we explore the effect of mergers on the morphology of radio-AGN. In §2, we explain the sample and data used in this work. In §3, we describe the methods used to create images of the radio-AGN and classify them. In §4, we present the results of our investigations into the morphology and number of radio-AGN in mergers versus non-mergers and compare them to optically identified AGN. In §5, we discuss the implications of our results and in §6 we summarize our findings. Throughout this paper we use a cosmology in which = 70 km s-1 Mpc-1, = 0.3, and =0.7.
@@ 2.1 The SDSS Galaxy Cluster Sample
The parent sample of galaxy clusters used for this study was obtained from Bilton et al. (2020). This galaxy cluster sample was assembled with Sloan Digital Sky Survey Data Release 8 galaxies (SDSS DR8; York et al. 2000; Aihara et al. 2011), which were constrained by parameters found in the literature, and compiled into the X-Ray Cluster Database (BAX; Sadat et al. 2004). Where the SDSS DR8 spectroscopic data is magnitude limited in the r-band to (Strauss et al., 2002). Furthermore, the MPA-JHU value-added catalogue was cross-matched to the DR8 membership, in particular for their stellar mass estimates, in order to maintain completeness (Kauffmann et al., 2003; Brinchmann et al., 2004; Salim et al., 2007).
@@
Abell 1367 | 11:44:29.5 | 19:50:20.6 | 0.022 | 58 | 0.245 | NM |
Abell 1656 | 12:59:48.7 | 27:58:50.5 | 0.023 | 96 | 0.117 | NM |
Abell 1795 | 13:49:00.5 | 26:35:06.8 | 0.062 | 46 | 0.127 | NM |
Abell 2029 | 15:10:56.0 | 05:44:41.0 | 0.077 | 98 | 0.307 | NM |
Abell 2061 | 15:21:15.3 | 30:39:16.7 | 0.078 | 95 | 0.0913 | NM |
Abell 2065 | 15:22:42.6 | 27:43:21.5 | 0.073 | 87 | 0.123 | NM |
Abell 2069 | 15:23:57.9 | 29:53:25.8 | 0.116 | 57 | 0.0765 | NM |
Abell 2199 | 16:28:38.5 | 39:33:06.0 | 0.030 | 120 | 0.16 | NM |
Abell 2255 | 17:12:31.0 | 64:05:33.3 | 0.081 | 83 | 0.18 | M |
Abell 1991 | 14:54:30.2 | 18:37:51.1 | 0.044 | 63 | 0.116 | M |
Abell 2033 | 15:11:28.1 | 06:21:43.7 | 0.082 | 41 | 0.184 | M |
Abell 426 | 03:18:36.4 | 41:30:54.2 | 0.018 | 64 | 0.269 | M |
@@
Note. — Sample of clusters used in this analysis. Cluster name, coordinates, redshift (), number of cluster members () identified by optical SDSS data in Bilton et al. (2020), global RMS value of LoTSS image (mJy beam-1), and dynamical state (), merging or non-merging, for clusters from Bilton et al. (2020) covered by LoTSS.
@@
Bilton et al. (2020) used the BAX galaxy cluster database to constrain the initial galaxy cluster sample to an X-ray luminosity range erg s-1. Applying these limits ensures that the most massive, well-assembled galaxy clusters are selected, while maintaining a sufficiently sized sample that can represent varying dynamical states within the finite limits of the -space that can be feasibly observed. After initial constraints were applied, the final cluster memberships for each cluster were then produced through surface caustics from the mass estimation methods of Diaferio & Geller (1997) and Diaferio (1999). These surface caustics vary as a function of projected radius from the cluster centre, thereby leading to the finalized membership being found within the confines of these caustics (see Gifford & Miller 2013; Gifford et al. 2013). Once the cluster galaxy membership had been ascertained, Bilton et al. (2020) proceeded to distinguish between relaxed and unrelaxed states, or ‘non-merging’ and ‘merging’ respectively. This was achieved via the incorporation of the Dressler & Shectman (1988) statistical test for substructure (also known as the -test), where the presence of substructure is used as a proxy for a galaxy cluster in a ‘merging’ or unrelaxed dynamical state. The details of this statistical test and its calculation are elaborated on in Bilton et al. (2020).
@@
Utilising the DR8 spectral lines for H and [NII] , so-called ‘WHAN’ diagrams can be used as a diagnostic to determine whether a cluster galaxy hosts an AGN or not (Cid Fernandes et al., 2010, 2011). Specifically, the equivalent width of H, EWHα, is compared to the logarithmic ratio of log10([NII]/H). The reduction of the number of emission lines for a WHAN diagram to two allows for mitigation against the SNR requirements for all emission lines, especially when compared to the Baldwin-Philips-Terlavich (BPT) diagnostic plot’s use of four emission lines (Baldwin et al., 1981). We refer to these WHAN selected AGN as optical AGN in the rest of this work. Furthermore, to reduce the number of interloping star-forming galaxies and Low Ionisation Emission Regions (LIERs), galaxies are classified as AGN if they possess emission ratios and strengths of log10([NII]/H) and EWHα 6Å respectively.
@@
During our initial analysis of the cluster member catalogs from Bilton et al. (2020), we discovered duplicate sources. We determined that these duplicates were a product of the fact that the SDSS DR8 spectroscopic catalog contains multiple measurements for some sources (along the plate overlap regions). We removed the 143 duplicate sources, which leaves us with a total of 2,298 objects.
@@
In addition to the duplicates found within each cluster, we also found 42 sources that were counted as members in both the Abell 2029 and Abell 2033 member catalogs. We note that Walker et al. (2012) show X-ray evidence of overlap between these two systems. To remedy this, for each of these 42 duplicate members we calculate a statistic, , which is based on both projected separation and relative velocity (Smith et al., 2004). The statistic is defined as
@@
where is the speed of light, is the redshift of the galaxy, is the cluster redshift, is the velocity dispersion of the cluster, is the projected radius of the galaxy from the cluster center in r200, and is the radius of the cluster in units of r200. We then assign the member to the cluster that minimizes (Smith et al., 2004). Minimizing , 37 of the duplicates are assigned to Abell 2029 and 5 assigned to Abell 2033.

### v3

# The Merging Galaxy Cluster Environment Affects the Morphology of Radio Active Galactic Nuclei

@@ Abstract
It has previously been found that the galaxy cluster environment can affect the fueling and evolution of active galactic nuclei (AGN). This work examines the effect of the merging cluster environment on the properties of radio AGN by comparing the radio morphology of cluster members in a sample of four merging and eight relaxed galaxy clusters at low redshift ( 0.2). Using 144 MHz data from the LOFAR Two-meter Sky Survey and Zooniverse, we classify the radio morphology of the radio-detected cluster members using the following morphology classes: compact, compact extended, extended, jetted, and disturbed. We find that the merging cluster environment has a statistically significant, higher population proportion of disturbed (bent and head-tail) sources, indicating that the merging environment can affect the morphology of cluster radio AGN. We also investigate the number of AGN that are detected in the radio data only and the number that are detected in both the radio and optical data in mergers and nonmergers. We find that the merging cluster environment has a higher population proportion of AGN that are identified only as radio AGN compared to AGN that are identified as both radio and optical AGN. Overall, we find that the merging environment affects certain radio AGN (disturbed and only radio-identified AGN), but not all.
@@ 1 Introduction
It is now well established that dense environments such as galaxy clusters affect the properties and evolution of the galaxies within them (see references detailed below). There are a variety of interactions between the galaxies, the intracluster medium (ICM; a diffuse, hot, ionized plasma that is gravitationally bound to the cluster), and the cluster potential that are possible and prevalent in clusters that affect properties and evolution of galaxies. Some examples of the possible interactions are galaxygalaxy harassment (repeated close encounters with neighbors; Moore et al. 1996, 1999), galaxy cluster potential tidal effects (Fujita, 1998; Natarajan et al., 1998), and galaxy interactions with the ICM, such as ram pressure stripping (Gunn & Gott, 1972) and strangulation (Larson et al., 1980). These interactions can affect the star formation rates (SFR; Kauffmann et al. (2004); Fassbender et al. (2014); Barsanti et al. (2018)), active galactic nuclei (AGN) activity (Galametz et al., 2009; Bufanda et al., 2017; Poggianti et al., 2017; Marshall et al., 2018; Mo et al., 2018), morphology (Dressler, 1980; Kauffmann et al., 2004; Sazonova et al., 2020), and gas supply of the cluster galaxies (Stroe et al., 2015; Poggianti et al., 2017; Jaffé et al., 2018; Cairns et al., 2019).
@@
One classic example of the effect of the cluster environment on its galaxies is the morphologydensity relation, in which, with increasing density, the fraction of elliptical galaxies increases and the fraction of spiral galaxies decreases such that clusters typically have an overabundance of elliptical galaxies compared to the field (Dressler, 1980; Kodama et al., 2001; Kauffmann et al., 2004; Sazonova et al., 2020). Coupled to the morphology-density relation is the SFR, where more star-formation is seen in galaxies at low densities and less at high densities (Kauffmann et al., 2004; Fassbender et al., 2014). There is an evolution of these properties as a function of redshift, in that, with decreasing redshift, the fraction of “red and dead” galaxies increases (De Lucia et al., 2004; Rudnick et al., 2009) and the fraction of star forming galaxies decreases (Margoniner et al., 2001; Poggianti et al., 2006; Saintonge et al., 2008). Another example of the environment affecting cluster galaxies is the case of “jellyfish” galaxies, which are galaxies typically located near the cores of the clusters that are experiencing ram pressure stripping, leaving a tail of gas behind them opposite to the direction of motion (Fumagalli et al., 2014; Ebeling et al., 2014; Jaffé et al., 2018).
@@
The cluster environment is also known to affect the properties and evolution of cluster galaxies that host AGN. AGN are powerful astrophysical sources that are fueled by accretion of matter onto the supermassive black hole at the center of the galaxy and emit characteristic radiation signatures across the full electromagnetic spectrum (Padovani et al., 2017). AGN can be identified in one or multiple wavelengths such as X-ray, optical, infrared, or radio. At all redshifts Galametz et al. (2009) and Mo et al. (2018) find that there is an excess of radio-identified AGN in clusters, specifically at the center of clusters compared to the field, which is more pronounced at high redshift (Galametz et al., 2009; Mo et al., 2020). Similarly, pertaining to X-ray, optical, and infrared identified AGN, the number density of AGN continually increases as a function of redshift in that there is a deficit of these types of AGN near the cluster center at (Pimbblet et al., 2013; Ehlert et al., 2014; Gordon et al., 2018), but at there is an excess of these types of AGN in clusters (Ruderman & Ebeling, 2005; Galametz et al., 2009; Fassbender et al., 2012; Martini et al., 2013; Alberts et al., 2016; Bufanda et al., 2017) which is most pronounced near the cluster center (Ruderman & Ebeling, 2005; Galametz et al., 2009; Fassbender et al., 2012; Alberts et al., 2016). The cluster environment seems to affect many different properties of radio AGN in particular, such as the size, luminosity, and morphology of the radio AGN (Prestage & Peacock, 1988; Gendre et al., 2013; Ineson et al., 2015; Croston et al., 2019; Garon et al., 2019; Moravec et al., 2020; Macconi et al., 2020; Shen et al., 2020; Morris et al., 2022). Further, bent-tail radio galaxies are a very well-known type of radio AGN that showcase the effect of the environment on radio AGN produced by the ram pressure of the dense environment (Miley et al., 1972; Begelman et al., 1979; O’Donoghue et al., 1993; Hardcastle et al., 2005; Morsony et al., 2013).
@@
Galaxy clusters undergo clustercluster mergers (henceforth mergers111In this work we refer to clustercluster mergers as “mergers,” which we differentiate from the typical galaxygalaxy mergers that many in the literature refer to as “mergers.”) as a result of hierarchical structure formation in the Universe (Voit, 2005; Springel et al., 2005; Boylan-Kolchin et al., 2009; Klypin et al., 2011; Pillepich et al., 2018; Planck Collaboration et al., 2020). Mergers have an immense and obvious impact on the properties of the ICM. For example, mergers cause the ICM distribution to be asymmetric as seen by X-ray (Buote, 2002) and SunyaevZel’dovich observations (Mroczkowski et al., 2019). Further, mergers drive intergalactic shocks and induce turbulence in the ICM (Sarazin, 2002; Markevitch, 2006; Basu et al., 2016). Lastly, mergers can create cold fronts in the ICM (Sarazin, 2002; Owers et al., 2009; Roediger et al., 2013) and disrupt the cool cores of clusters (Sarazin, 2002).
@@
Although the effect of mergers on the ICM is clear, the impact on the cluster galaxies is less so. There is contradictory evidence that mergers can either increase (Owen et al., 1999; Miller & Owen, 2003; Hwang & Lee, 2009) or decrease (Hwang & Lee, 2009; Mansheim et al., 2017) star formation in cluster galaxies. One study found that a dynamically younger cluster had a higher fraction of galaxies with a high SFR (0.17 yr-1, Stroe et al. 2015). There is also evidence that a brief period of enhanced, triggered star formation may be followed by subsequent quenching (Stroe et al., 2015; Cairns et al., 2019). These results are attributed to a disturbed ICM affecting the molecular gas reservoir of the cluster galaxies, with the disturbance exciting star formation and leading to a rapid consumption of the molecular gas (Stroe et al., 2015; Cairns et al., 2019).
@@
Recent work has found that the AGN population is impacted by clustercluster mergers as well. The radio AGN population in particular has been studied in depth in individual merging systems and found to have a higher frequency compared to the radio AGN frequency in more relaxed systems (e.g., Miller & Owen, 2003; Moravec et al., 2020). Similarly, Noordeh et al. (2020) found enhanced X-ray AGN activity in the most dynamically disturbed of the seven clusters they studied. Bilton et al. (2020) find that mergers hold relatively kinematically younger AGN subpopulations that have recently coalesced into a common potential (similar to a first infall population of galaxies) compared to relaxed clusters. In general, Miller & Owen (2003); Hwang & Lee (2009) and Sobral et al. (2015) find that mergers can provoke AGN activity. However, the physical mechanism by which mergers instigate AGN activity and its effects on radio AGN properties are unknown.
@@
In this paper, we explore the effect of mergers on the morphology of radio AGN. In §2, we explain the sample and data used in this work. In §3, we describe the methods used to create images of the radio AGN and classify them. In §4, we present the results of our investigations into the morphology and number of radio AGN in mergers versus nonmergers and compare them to optically identified AGN. In §5, we discuss the implications of our results. In §6 we summarize our findings. Throughout this paper we use a cosmology in which 70 kms-1 Mpc-1, 0.3, and 0.7.
@@ 2.1 The SDSS Galaxy Cluster Sample
The parent sample of galaxy clusters used for this study was obtained from Bilton et al. (2020). This galaxy cluster sample was assembled with Sloan Digital Sky Survey Data Release 8 galaxies (SDSS DR8; York et al. 2000; Aihara et al. 2011), which were constrained by parameters found in the literature, and compiled into the X-Ray Cluster Database (BAX; Sadat et al. 2004). The SDSS DR8 spectroscopic data are magnitude limited in the band to (Strauss et al., 2002). Furthermore, the MPA-JHU value-added catalog was cross-matched to the DR8 membership, in particular for their stellar mass estimates, in order to maintain completeness (Kauffmann et al., 2003; Brinchmann et al., 2004; Salim et al., 2007).
@@
A1367 | 11:44:29.5 | 19:50:20.6 | 0.022 | 58 | 0.245 | NM |
A1656 | 12:59:48.7 | 27:58:50.5 | 0.023 | 96 | 0.117 | NM |
A1795 | 13:49:00.5 | 26:35:06.8 | 0.062 | 46 | 0.127 | NM |
A2029 | 15:10:56.0 | 05:44:41.0 | 0.077 | 98 | 0.307 | NM |
A2061 | 15:21:15.3 | 30:39:16.7 | 0.078 | 95 | 0.0913 | NM |
A2065 | 15:22:42.6 | 27:43:21.5 | 0.073 | 87 | 0.123 | NM |
A2069 | 15:23:57.9 | 29:53:25.8 | 0.116 | 57 | 0.0765 | NM |
A2199 | 16:28:38.5 | 39:33:06.0 | 0.030 | 120 | 0.16 | NM |
A2255 | 17:12:31.0 | 64:05:33.3 | 0.081 | 83 | 0.18 | M |
A1991 | 14:54:30.2 | 18:37:51.1 | 0.044 | 63 | 0.116 | M |
A2033 | 15:11:28.1 | 06:21:43.7 | 0.082 | 41 | 0.184 | M |
A426 | 03:18:36.4 | 41:30:54.2 | 0.018 | 64 | 0.269 | M |
@@
Note. — Sample of clusters used in this analysis. Columns provide cluster name, coordinates, redshift (), number of cluster members () identified by optical SDSS data in Bilton et al. (2020), global RMS value of LoTSS image (mJy beam-1), and dynamical state (, merging or nonmerging) for clusters from Bilton et al. (2020) covered by LoTSS.
@@
Bilton et al. (2020) used the BAX galaxy cluster database to constrain the initial galaxy cluster sample to an X-ray luminosity range erg s-1 erg s-1. Applying these limits ensures that the most massive, well-assembled galaxy clusters are selected, while maintaining a sufficiently sized sample that can represent varying dynamical states within the finite limits of the -space that can be feasibly observed. After initial constraints were applied, the final cluster memberships for each cluster were then produced through surface caustics from the mass estimation methods of Diaferio & Geller (1997) and Diaferio (1999). These surface caustics vary as a function of projected radius from the cluster centre, thereby leading to the finalized membership being found within the confines of these caustics (see Gifford & Miller 2013; Gifford et al. 2013). Once the cluster galaxy membership had been ascertained, Bilton et al. (2020) proceeded to distinguish between relaxed and unrelaxed states, or ‘nonmerging’ and ‘merging,’ respectively. This was achieved via the incorporation of the Dressler & Shectman (1988) statistical test for substructure (also known as the -test), where the presence of substructure is used as a proxy for a galaxy cluster in a ‘merging’ or unrelaxed dynamical state. The details of this statistical test and its calculation are elaborated on in Bilton et al. (2020).
@@
Utilizing the DR8 spectral lines for H and [N] , so-called ‘WHAN’ diagrams can be used as a diagnostic to determine whether a cluster galaxy hosts an AGN or not (Cid Fernandes et al., 2010, 2011). Specifically, the equivalent width of H, EWHα, is compared to the logarithmic ratio of log10([N ii]/H). The reduction of the number of emission lines for a WHAN diagram to two allows for mitigation against the signal-to-noise (S/N) requirements for all emission lines, especially when compared to the BaldwinPhilipsTerlavich (BPT) diagnostic plot’s use of four emission lines (Baldwin et al., 1981). We refer to these WHAN-selected AGN as optical AGN in the rest of this work. Furthermore, to reduce the number of interloping star-forming galaxies and low ionisation emission regions, galaxies are classified as AGN if they possess emission ratios and strengths of log10([N ii]/H) and EWHα 6Å respectively.
@@
During our initial analysis of the cluster member catalogs from Bilton et al. (2020), we discovered duplicate sources. We determined that these duplicates were a product of the fact that the SDSS DR8 spectroscopic catalog contains multiple measurements for some sources (along the plate overlap regions). We removed the 143 duplicate sources, which leaves us with a total of 2298 objects.
@@
In addition to the duplicates found within each cluster, we also found 42 sources that were counted as members in both the A2029 and A2033 member catalogs. We note that Walker et al. (2012) show X-ray evidence of overlap between these two systems. To remedy this, for each of these 42 duplicate members we calculate a statistic, , which is based on both projected separation and relative velocity (Smith et al., 2004). The statistic is defined as
@@
where is the speed of light, is the redshift of the galaxy, is the cluster redshift, is the velocity dispersion of the cluster, is the projected radius of the galaxy from the cluster center in r200, and is the radius of the cluster in units of r200. We then assign the member to the cluster that minimizes (Smith et al., 2004). Minimizing , 37 of the duplicates are assigned to A2029 and 5 are assigned to A2033.